OpenStudy (anonymous):

A game involves tossing two coins. A player wins $1.00 if both tosses result in heads. What should you pay to play this game in order to make it a fair game? Explain your answer.

OpenStudy (anonymous):

y = the amount to pay Coin Winnings Probability Expectation tosses or losses of this per game X P(X) X•P(X) --------------------------------------------- HH $1-y ¼ (1-y)(¼) HT -y ¼ -y(¼) TH -y ¼ -y(¼) TT -y ¼ -y(¼) Sum of expectations = 0 (1-y)(¼) - y(¼) - y(¼) -y(¼) = 0 Multiply through by 4 (1-y) - y - y - y = 0 1 - y - y - y - y = 0 1 - 4y = 0 -4y = -1 y = ¼ = .25 = 25 cents This is what I came up with but I don't understand why the sum is 0

OpenStudy (anonymous):

there are four different results that can happen. Having both coins tossed as heads is one of the four results that might occur. so there is a one in four chance that you can win $1. Divide $1 by four and you get $0.25, which would make it a fair game.

OpenStudy (anonymous):

Cool much more simple then the way I did it - Thanks